OpenStudy (kewlgeek555):

Describe the solution(s) to the equation x^2 + 5x + 4 = 0 by just determining discriminant. Show your work. Huh? cx

OpenStudy (solomonzelman):

Do you know what a discriminant is? If so, then you also know that the general fom of this type of equations is ax^2+bx+c=0 Can you plug in your numbers ?

OpenStudy (kewlgeek555):

I know about the general form of the equation is, but I seemed to have forgotten what the discriminant is. :3

OpenStudy (solomonzelman):

The discriminant is b^2-4ac (the thing inside the square rot in the quadratic formula)

OpenStudy (kewlgeek555):

Is the discriminant just \[b^2-4ac\]or is it the actual \[\sqrt{b^2-4ac}\]?

OpenStudy (johnweldon1993):

Square root is not included...so the first one you posted @kewlgeek555

OpenStudy (kewlgeek555):

Okay, thanks. So I have to plug in the values, right? That would get me... \[5^s-4(1)(4)\]Right?

OpenStudy (kewlgeek555):

Oops, I meant to put five to the power of two, not eight. cx

OpenStudy (johnweldon1993):

Now...the question doesn't state to SAY the answers to the question...merely to "describe" them The objective of the discriminant is to tell you how many, and of what type, of solutions there are to your polynomial

OpenStudy (johnweldon1993):

Lol yeah fix that exponent and then you will be correct... \[\large 5^2 - 4(1)(4) = \space ?\]

OpenStudy (kewlgeek555):

Oh okay. Thanks! Is there another method other than the discriminant to solve equations like these? (ax^2 + bx + c = 0)? Are there specific times when you need to us them?

OpenStudy (johnweldon1993):

Not done yet :P still need to solve the \[\large 5^2 - 4(1)(4) = \space ?\]

OpenStudy (johnweldon1993):

When you solve that...look at your result... The results of the discriminant will show you how many, and of what type, of solutions you have for this polynomial... If the result > 0 you will have 2 real roots If the result = 0 you will have 1 real root If the result < 0 you will have 2 complex roots

OpenStudy (johnweldon1993):

Also note...the discriminant does NOT give you the solutions to the equation...just tells you about them

OpenStudy (kewlgeek555):

\[5^2-4(1)(4)\]\[25-4(1)(4)\]\[25-4(4)\]\[25-16\]\[9\]Okay, so since the result is greater than zero, I have two real roots. c:

OpenStudy (johnweldon1993):

That is correct...2 real roots That will be your "description" of your solutions

OpenStudy (johnweldon1993):

Now...to answer your other question... is there a way to solve a polynomial? (ax^2 + bx + c) Yes of course...the quadratic equation *this WILL give you your solutions* \[\large \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\]

OpenStudy (kewlgeek555):

Okay, and is there another way? Because it is asking me for two methods. Isn't there the b over 2a method, or is that for something else?

OpenStudy (johnweldon1993):

No, that would solve for the vertex of your parabola... 2 methods huh? well...you can use that equation Or....you can factor a.k.a make ax^2 + bx + c look something like (x + a)(x + b) *note...the 'a' in the first equation does NOT = the 'a' in the second part* it is just the form I am describing

OpenStudy (kewlgeek555):

Oh okay, I understand. :D

OpenStudy (kewlgeek555):

Thanks, nice to see you again! ;D

OpenStudy (johnweldon1993):

Anytime!